A man cycles to his destination with the wind at 8 m/s , and returns home against the wind at 4 m/s . What is his average speed?
1
Expert's answer
2020-10-05T10:55:30-0400

"v=\frac{s+s}{s/8+s/4}=\frac{2}{1/8+1/4}=5.33(m/s)"